Hexane (C6H14) is used as an organic solvent according to US EPA, there is no information its carcinogenic
effects in humans or animals. If it will be treated as solute of its own, what is the molarity of hexane?
Density= 0.655 g/mL
1
Expert's answer
2020-10-22T14:18:52-0400

Given,

Density of the hexane,D = 0.655 g/ml

As we know,

D = (M/V)

Or, M = V×D


So,

mass of (1L=) 1000 ml Hexane

= (1000/0.655) g

= 1526.717 g


Molar mass of Hexane (C6H14)

= (12×6+1×12) g/mol

= 84 g/mol


1L = 1000 ml hexane will have the mass = 1526.717 g

=[ 1526.717 g / 84( g/mol ) ]

= 18.175 mol


As, 1 L of volume contain 18.175 mol of Hexane.


So,

If Hexane is treated as solute of its own , the Molarity of Hexane = 18.175 mol /L